OpenStudy (anonymous):

Given log v2 x = 3, find log v2 (4x) and log v2 x^5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

v2?

OpenStudy (anonymous):

how can I type log2(x)... just like that?

OpenStudy (anonymous):

as it log base 2, not log base 10

OpenStudy (anonymous):

x to base 2

OpenStudy (anonymous):

k the first answer is 5

OpenStudy (anonymous):

second answer is 15

OpenStudy (anonymous):

log 4x = log 4 + logx = 2 + 3 = 5

OpenStudy (anonymous):

log (x^5) = 5logx = 15

OpenStudy (anonymous):

got it?>

OpenStudy (anonymous):

yeah, I see how to do it, thank you :) (I'm new, how do I update the "needs a good answer" status and thank you for your help? :) )

OpenStudy (anonymous):

theres a button in blue to the right..."Good Answer" where ive given any answer of mine

OpenStudy (anonymous):

Thank you him1618.